Check whether the limit of the function 6 2
3
2
3
( , )
x y
x y
f x y
+
= exists as )0,0(
1
Expert's answer
2020-05-07T19:19:10-0400

Consider


"\lim\limits_{(x,y)\to(0,0)}{6x^2y^3 \over x^2+y^3}"

Suppose "\varepsilon>0"


"\bigg|{6x^2y^3 \over x^2+y^3}\bigg|=6\bigg|{y^3 \over x^2+y^3}\bigg|x^2"

Note that


"\bigg|{y^3 \over x^2+y^3}\bigg|\leq1, x^2\leq x^2+y^2=\delta^2"

So


"6\bigg|{y^3 \over x^2+y^3}\bigg|x^2<6\cdot1\cdot\delta^2"

If we choose "\delta=\sqrt{\dfrac{\varepsilon}{6}}" then


"6\bigg|{y^3 \over x^2+y^3}\bigg|x^2<6\cdot1\cdot\delta^2=\varepsilon"

"|f(x,y)-0|<\varepsilon, if\ \sqrt{x^2+y^2}<\delta"

Hence


"\lim\limits_{(x,y)\to(0,0)}{6x^2y^3 \over x^2+y^3}=0"